OpenStudy (anonymous):

solve the variable for : 3w = 4 1/5

OpenStudy (anonymous):

\[3w = 4 \frac{1}{5}\]

OpenStudy (anonymous):

\[4\frac{1}{5}=\frac{21}{5}\]

OpenStudy (anonymous):

replace it and solve for w

OpenStudy (anonymous):

Replace w with 21/5?

OpenStudy (anonymous):

what now? if you replace w then we have no w to solve for

OpenStudy (anonymous):

\[3w=\frac{21}{5}\]

OpenStudy (anonymous):

7/5

OpenStudy (anonymous):

nice that is right!

OpenStudy (anonymous):

tysm!